OpenStudy (anonymous):

what is the distance between the points (8,8) and (3,20)

OpenStudy (anonymous):

apply your distance forumula. d = sqrt [( x2-x1)^2+(y2-y1)^2] so you get ... d =sqrt [ (3-8)^2 + (20-8)^2 ] = sqrt [ (-5)^2 + (12)^2 ] = sqrt [ 25 + 144 ] = sqrt [169 ] = 13 your welcome.

OpenStudy (anonymous):

thank youuu so much

OpenStudy (anonymous):

no problemo